OpenStudy (anonymous):

please help me find the 2 asymptotes of y=2sec(x/4). I know that the period is 8pi but i don't know how to find the asymptotes

OpenStudy (sidsiddhartha):

y=2sec(x/4)=2/cos(x/4) equate cos(x/4)=0 x/4=(2n+1)*(pi/2) now u can do it

OpenStudy (anonymous):

to find the asymptotes after that would i put 0 and 1 in place of n? @sidsiddhartha

OpenStudy (sidsiddhartha):

that'll do

OpenStudy (anonymous):

okay thanks, once i get to that point and have the formula its pretty easy, but i have two more problems similar to this one and I'm not sure how to get to the point of finding the formula, could you help me understand that better?

OpenStudy (anonymous):

by the way for the 2 asymptotes i got 2pi and 6pi does that seem right? :o

OpenStudy (sidsiddhartha):

yes read this every things easily explained here http://www.purplemath.com/modules/triggrph2.htm

OpenStudy (anonymous):

okay, thank you! :)

OpenStudy (sidsiddhartha):

welcome